# Testing the Equivalence Principle on Cosmological Scales Using Peculiar Acceleration Power Spectrum

### v2

@@ I Introduction
The celebrated general theory of relativity is undoubtedly founded upon the Equivalence Principle (EP), which represents one of its cornerstones. EP states that all non-gravitational phenomena are locally unaffected by gravity, provided that the experiments are performed in a freely falling frame. A specific consequence of this principle is that all objects (including light) fall in accordance with the same laws. Thus far, EP has been subjected to rigorous testing within the solar system (Will, 2014; Touboul et al., 2017). However, its applicability on cosmological scales when considering dark matter (DM) and dark energy remains uncertain. In essence, the large-scale structure of the Universe contains valuable information about EP, which can be used to directly evaluate its validity. To date, only indirect methods have been proposed to test EP (Kesden and Kamionkowski, 2006a, b; Kehagias et al., 2014; Creminelli et al., 2014; Bonvin and Fleury, 2018; Umeh et al., 2021). In this study, we propose a direct test of (weak) EP via the density-weighted peculiar acceleration power spectrum, which can be quantified through the redshift drift technique.
@@
The measurement of redshift drift plays an essential role in real-time cosmology, as it is equivalent to observing the expansion of the Universe in real time. Moreover, the redshift drift is a model-independent method of probing the expansion of the Universe, thereby eliminating assumptions about geometry, clustering, or gravitational behavior. One of the key challenges in measuring redshift drifts arises from the vast difference between the time scales relevant to cosmology and human observation, requiring extremely precise measurements. The primary sources of uncertainty in redshift drift measurements arise from shot noise and redshift measurement precision. Addressing these challenges requires both a high number density of galaxies and spectroscopy with ultra-high spectral resolution.
@@ II.1 Estimator
where, in the second equality, the constant term $a_{\rm cos}$ contributes only to the $k=0$ mode and is therefore ignored. This is because the $k=0$ mode represents the spatially uniform component of the field, which does not carry information about the spatial structure or clustering of galaxies. In the last equality, we restrict our analysis to linear scales and neglect the second-order term $\left\langle\delta_{g}a_{\rm pec,\|}\right\rangle$ where the galaxy density fluctuation $\delta_{g}$ is much smaller than unity. In Appendix A, we present the theoretical formulation of the linear $P_{a_{\parallel}a_{\parallel}}$ and compare it with simulation measurements to test its validity  [^1] How the violation of this linear approximation will effect the accuracy of the estimator is tested in Sect. III, where we demonstrate that even its breakdown due to non-linearity on small scales does not significantly affect our results..
@@ II.2 Simulation
Ideally, the test should be performed on a simulation combining sufficiently large volume and high mass resolution to both resolve small-mass halos and reproduce the tracer number density expected in upcoming HI surveys like SKA. Due to the limited volume and resolution of our current simulation, we instead employ a two-pronged validation approach: one test utilizes DM particles as tracers to minimize shot noise and isolate the impact of redshift measurement uncertainty on the $E_{\rm ep}$ detection, while the other uses DM halos to assess how shot noise contributes to the statistical uncertainty in the $E_{\rm ep}$ measurement, and how the non-linear evolution of structure formation may introduce systematic bias in the interpretation of $E_{\rm ep}$
@@ II.2.1 DM particle mock generation
To mimic realistic observational uncertainties, we simulate redshift measurement errors by adding independent Gaussian noise to the redshift of each particle in both datasets. The noise was generated with a mean of zero and a standard deviation of $\sigma_{z_{\parallel}}=f\sigma_{\rm SKA}/t_{\rm obs}$, where $\sigma_{\rm SKA}=10~{\rm m/s}$ (Klöckner et al., 2015) represents the typical redshift measurement uncertainty expected from SKA observations (an integration time with $\mathcal{O}(1)-\mathcal{O}(10)$ hours per pointing), and $t_{\rm obs}=12$ years. Here the parameter $f$ is a scaling factor applied to $\sigma_{\rm SKA}$. In this study, we consider two noise levels to assess their impact on SKA observation accuracy: a low-noise case with $f=0.001$, and a high-noise case with $f=0.002$. The latter corresponds to a doubled noise amplitude, resulting in four times higher noise power in the power spectrum.
@@ II.2.2 DM Halo Mock generation
To assess the sensitivity of our estimator to violations of EP, we introduce artificial EP-breaking effects into the halo catalogs as follows. For each halo of mass $M$, we rescale its line-of-sight (LOS) acceleration by a mass-dependent factor $\lambda$:
@@
This model is not derived from any specific underlying theory, but instead serves as a controlled, parameterized prescription to introduce EP violations as a function of halo mass. It is a purely phenomenological framework, designed to systematically probe the sensitivity of the estimator to such violations. When $\alpha=0$, the EP is exactly preserved. Nonzero values of $\alpha$ introduce controlled deviations from EP. In this study, the characteristic mass scale is fixed at $M_{*}=10^{12}M_{\odot}$, as it represents the characteristic mass of typical galaxies.
We fix $\alpha=0.1$ and vary $\beta$ to control the mass dependence of the EP violation. Within this phenomenological framework, we consider representative values of $\beta=0.1$, $0.3$, and $0.5$ to assess how different scalings of the EP-breaking term influence the response of our estimator. Under these parameter choices, the relative deviations of $a_{||}$ from $\lambda$ in the small-mass group reach approximately 10-15% for $\beta=0.1$ and $0.3$, and up to 18% for $\beta=0.5$. In contrast, for the large-mass group, the deviations are more significant–about 20% for $\beta=0.1$, 100% for $\beta=0.3$, and a factor of 3.9–8.1 for $\beta=0.5$.
@@ III Results
To quantitatively characterize the degree of deviation from the expected value, we propose a simple one-parameter model to fit the observed data $E^{\rm obs}_{\rm ep}(k)$ over $k$ bins. The parameterization is expressed as:
@@ III.2 Results from DM Halo Mock
The results presented in Tab. 3 quantitatively illustrate the sensitivity of the $\delta_{\rm ep}$ estimator to different levels of EP violation, parameterized by $(\alpha,\beta)$. For the no-violation case with $\alpha=0$ and $\beta=0$, the best-fit $\delta_{\rm ep}$ values remain close to zero across all redshifts, ranging from approximately 0.08-0.08 to $0.05$, and are consistent with the $1\sigma$ uncertainties (about 0.040.060.04-0.06). Corresponding $\chi^{2}$ values range from 4.38 to 11.63, indicating statistically acceptable fits.
@@ IV Conclusion
The (weak) Equivalence Principle (EP) has been extensively tested within the Solar System; however, its validity on cosmological scales remains an open question. In this study, we propose a novel approach to test EP on cosmological scales by measuring the peculiar acceleration power spectrum of galaxies through the redshift drift technique.
To achieve this, we develop an EP estimator, $E_{\rm ep}$, designed to evaluate the consistency of peculiar acceleration power spectra across different tracers. If EP holds on cosmological scales for these tracers, the ensemble average of $E_{\rm ep}$ is expected to equal one. As a first step toward testing EP on cosmological scales, we validate this estimator using N-body simulations. Applying this approach, we measure $E_{\rm ep}$ at four redshift bins ($z=[0,0.5,1,1.5]$) on the linear scales of $k\in[0.007,0.2]~h/\rm Mpc$. This analysis is conducted using two distinct tests: i) DM particle mock data, incorporating redshift measurement errors, and ii) mock data for small-mass and large-mass DM halos.
@@
Acknowledgement: We thank Volker Springel for useful discussions. This work is supported by National SKA Program of China (2020SKA0110401, 2020SKA0110402, 2020SKA0110100), the National Key R&D Program of China (2020YFC2201600, 2018YFA0404504, 2018YFA0404601), the National Science Foundation of China (12203107, 12473097, 12073088, 12373005, 12273121), the China Manned Space Project with No. CMS-CSST-2021 (A02, A03, B01), the Guangdong Basic and Applied Basic Research Foundation (2024A1515012309), and the Fundamental Research Funds for the Central Universities, Sun Yat-sen University (No. 24qnpy122). We also wish to acknowledge the Beijing Super Cloud Center (BSCC) and Beijing Beilong Super Cloud Computing Co., Ltd (http://www.blsc.cn/) for providing HPC resources that have significantly contributed to the research results presented in this paper.
@@ Appendix A Validation of acceleration power spectrum
In Fig. 3, we compare the power spectrum $P_{a_{\|}a_{\|}}$ with the corresponding measurements derived from low- and high-mass halo catalogs at four distinct redshifts. The shot noise contribution has been subtracted from the measured spectra. On linear scales ($k\lesssim 0.06~h/\rm Mpc$), the measurements exhibit excellent agreement with the predictions of linear theory for both halo mass bins. This agreement supports the earlier argument in the text that higher-order terms, such as $\delta_{g}a_{\mathrm{pec},\|}$, can be neglected on linear scales. At smaller scales, however, the power spectra deviate from the linear predictions due to the influence of non-linear evolution and halo assembly processes.

### v3

@@ I Introduction
The celebrated general theory of relativity is undoubtedly founded upon the equivalence principle (EP), which represents one of its cornerstones. EP states that all nongravitational phenomena are locally unaffected by gravity, provided that the experiments are performed in a freely falling frame. A specific consequence of this principle is that all objects (including light) fall in accordance with the same laws. Thus far, EP has been subjected to rigorous testing within the solar system (Will, 2014; Touboul et al., 2017). However, its applicability on cosmological scales when considering dark matter (DM) and dark energy remains uncertain. In essence, the large-scale structure of the Universe contains valuable information about EP, which can be used to directly evaluate its validity. To date, only indirect methods have been proposed to test EP (Kesden and Kamionkowski, 2006a, b; Kehagias et al., 2014; Creminelli et al., 2014; Bonvin and Fleury, 2018; Umeh et al., 2021). In this study, we propose a direct test of (weak) EP via the density-weighted peculiar acceleration power spectrum, which can be quantified through the redshift drift technique.
@@
The measurement of redshift drift plays an essential role in real-time cosmology, as it is equivalent to observing the expansion of the Universe in real time. Moreover, the redshift drift is a model-independent method of probing the expansion of the Universe, thereby eliminating assumptions about geometry, clustering, or gravitational behavior. One of the key challenges in measuring redshift drifts arises from the vast difference between the time scales relevant to cosmology and human observation, requiring extremely precise measurements. The primary sources of uncertainty in redshift drift measurements arise from shot noise and redshift measurement precision. Addressing these challenges requires both a high number density of galaxies and spectroscopy with ultrahigh spectral resolution.
@@ II.1 Estimator
where, in the second equality, the constant term $a_{\rm cos}$ contributes only to the $k=0$ mode and is therefore ignored. This is because the $k=0$ mode represents the spatially uniform component of the field, which does not carry information about the spatial structure or clustering of galaxies. In the last equality, we restrict our analysis to linear scales and neglect the second-order term $\left\langle\delta_{g}a_{\rm pec,\|}\right\rangle$ where the galaxy density fluctuation $\delta_{g}$ is much smaller than unity. In Appendix A, we present the theoretical formulation of the linear $P_{a_{\parallel}a_{\parallel}}$ and compare it with simulation measurements to test its validity  [^1] How the violation of this linear approximation will effect the accuracy of the estimator is tested in Sect. III, where we demonstrate that even its breakdown due to nonlinearity on small scales does not significantly affect our results..
@@ II.2 Simulation
Ideally, the test should be performed on a simulation combining sufficiently large volume and high mass resolution to both resolve small-mass halos and reproduce the tracer number density expected in upcoming HI surveys like SKA. Due to the limited volume and resolution of our current simulation, we instead employ a two-pronged validation approach: one test utilizes DM particles as tracers to minimize shot noise and isolate the impact of redshift measurement uncertainty on the $E_{\rm ep}$ detection, while the other uses DM halos to assess how shot noise contributes to the statistical uncertainty in the $E_{\rm ep}$ measurement, and how the nonlinear evolution of structure formation may introduce systematic bias in the interpretation of $E_{\rm ep}$
@@ II.2.1 DM particle mock generation
To mimic realistic observational uncertainties, we simulate redshift measurement errors by adding independent Gaussian noise to the redshift of each particle in both datasets. The noise was generated with a mean of zero and a standard deviation of $\sigma_{z_{\parallel}}=f\sigma_{\rm SKA}/\sqrt{t_{\rm obs}}$, where $\sigma_{\rm SKA}=10~{\rm m/s}$ (Klöckner et al., 2015) represents the typical redshift measurement uncertainty expected from SKA observations (an integration time with $\mathcal{O}(1)-\mathcal{O}(10)$ hours per pointing), and $t_{\rm obs}=12$ years. Here the parameter $f$ is a scaling factor applied to $\sigma_{\rm SKA}$. In this study, we consider two noise levels to assess their impact on SKA observation accuracy: a low-noise case with $f=0.001$, and a high-noise case with $f=0.002$. The latter corresponds to a doubled noise amplitude, resulting in four times higher noise power in the power spectrum.
@@ II.2.2 DM Halo Mock generation
To assess the sensitivity of our estimator to violations of EP, we introduce artificial EP-breaking effects into the halo catalogs as follows. For each halo of mass $M$, we rescale its LOS acceleration by a mass-dependent factor $\lambda$:
@@
This model is not derived from any specific underlying theory, but instead serves as a controlled, parametrized prescription to introduce EP violations as a function of halo mass. It is a purely phenomenological framework, designed to systematically probe the sensitivity of the estimator to such violations. When $\alpha=0$, the EP is exactly preserved. Nonzero values of $\alpha$ introduce controlled deviations from EP. In this study, the characteristic mass scale is fixed at $M_{*}=10^{12}M_{\odot}$, as it represents the characteristic mass of typical galaxies.
We fix $\alpha=0.1$ and vary $\beta$ to control the mass dependence of the EP violation. Within this phenomenological framework, we consider representative values of $\beta=0.1$, $0.3$, and $0.5$ to assess how different scalings of the EP-breaking term influence the response of our estimator. Under these parameter choices, the relative deviations of $a_{||}$ from $\lambda$ in the small-mass group reach approximately 10%-15% for $\beta=0.1$ and $0.3$, and up to 18% for $\beta=0.5$. In contrast, for the large-mass group, the deviations are more significant–about 20% for $\beta=0.1$, 100% for $\beta=0.3$, and a factor of 3.9–8.1 for $\beta=0.5$.
@@ III Results
To quantitatively characterize the degree of deviation from the expected value, we propose a simple one-parameter model to fit the observed data $E^{\rm obs}_{\rm ep}(k)$ over $k$ bins. The parametrization is expressed as:
@@ III.2 Results from DM Halo Mock
The results presented in Tab. 3 quantitatively illustrate the sensitivity of the $\delta_{\rm ep}$ estimator to different levels of EP violation, parametrized by $(\alpha,\beta)$. For the no-violation case with $\alpha=0$ and $\beta=0$, the best-fit $\delta_{\rm ep}$ values remain close to zero across all redshifts, ranging from approximately 0.08-0.08 to $0.05$, and are consistent with the $1\sigma$ uncertainties (about 0.040.060.04-0.06). Corresponding $\chi^{2}$ values range from 4.38 to 11.63, indicating statistically acceptable fits.
@@ IV Conclusion
The (weak) EP has been extensively tested within the Solar System; however, its validity on cosmological scales remains an open question. In this study, we propose a novel approach to test EP on cosmological scales by measuring the peculiar acceleration power spectrum of galaxies through the redshift drift technique.
To achieve this, we develop an EP estimator, $E_{\rm ep}$, designed to evaluate the consistency of peculiar acceleration power spectra across different tracers. If EP holds on cosmological scales for these tracers, the ensemble average of $E_{\rm ep}$ is expected to equal 1. As a first step toward testing EP on cosmological scales, we validate this estimator using N-body simulations. Applying this approach, we measure $E_{\rm ep}$ at four redshift bins ($z=[0,0.5,1,1.5]$) on the linear scales of $k\in[0.007,0.2]~h/\rm Mpc$. This analysis is conducted using two distinct tests: i) DM particle mock data, incorporating redshift measurement errors, and ii) mock data for small-mass and large-mass DM halos.
@@
Acknowledgments: We thank Volker Springel for useful discussions. We thank the anonymous referee for useful suggestions that significantly improve the quality of paper. This work is supported by the National SKA Program of China (2020SKA0110401, 2020SKA0110402, 2020SKA0110100), the National Key R&D Program of China (2020YFC2201600, 2018YFA0404504, 2018YFA0404601), the National Science Foundation of China (12203107, 12473097, 12073088, 12373005, 12273121), the China Manned Space Project with No. CMS-CSST-2021 (A02, A03, B01), the Guangdong Basic and Applied Basic Research Foundation (2024A1515012309), and the Fundamental Research Funds for the Central Universities, Sun Yat-sen University (No. 24qnpy122). We also wish to acknowledge the Beijing Super Cloud Center (BSCC) and Beijing Beilong Super Cloud Computing Co., Ltd (http://www.blsc.cn/) for providing HPC resources that have significantly contributed to the research results presented in this paper.
@@ Appendix A Validation of acceleration power spectrum
In Fig. 3, we compare the power spectrum $P_{a_{\|}a_{\|}}$ with the corresponding measurements derived from low- and high-mass halo catalogs at four distinct redshifts. The shot noise contribution has been subtracted from the measured spectra. On linear scales ($k\lesssim 0.06~h/\rm Mpc$), the measurements exhibit excellent agreement with the predictions of linear theory for both halo mass bins. This agreement supports the earlier argument in the text that higher-order terms, such as $\delta_{g}a_{\mathrm{pec},\|}$, can be neglected on linear scales. At smaller scales, however, the power spectra deviate from the linear predictions due to the influence of nonlinear evolution and halo assembly processes.